OpenStudy (anonymous):

32(3x^4-8x^2-16)= 0 Is 2nd derivative so this will give me my inflection points. help please

OpenStudy (science0229):

If that is your 2nd derivative, first solve the equation for all real numbers.

OpenStudy (anonymous):

thats what I need help with

OpenStudy (anonymous):

let u = x^2 to convetr this to a quadratic in u, then solve using standard solution for roots of quadratic.

OpenStudy (anonymous):

so I will have 3U^2-8U-16?

OpenStudy (anonymous):

I get 4, -1.3333333.. does this mean x= sqrt4, -sqrt1.3333333?

OpenStudy (anonymous):

That's the conversion. You have solved for U but U is x^2, so take the square root of the solutions for U to get x.

OpenStudy (anonymous):

look at the graph and see if this makes sense.

OpenStudy (anonymous):

doesnt look right on the graph at all

OpenStudy (science0229):

Quick tip: convert your original U to fraction.

OpenStudy (anonymous):

huh?

OpenStudy (science0229):

3U^2-8U-16=0 Factor to get (U-4)(3U-4)=0 U=4 or U=4/3 So, x=2, sqrt(4/3), -2, -sqrt(4/3)

OpenStudy (anonymous):

i should have just factored and ignored the quadratic on this one,right?

OpenStudy (science0229):

yes.

OpenStudy (science0229):

Do you know what to do from here on?

OpenStudy (anonymous):

let me give you the entire 2nd derivative. all of that over (x^2-4)^2

OpenStudy (anonymous):

actually, its 32(3x^4-3x^2-16)/(x^2-4)^2

OpenStudy (science0229):

Oh, then it gets easier.

OpenStudy (anonymous):

-8x^2 not 3

OpenStudy (science0229):

Because the denominator can't be negative, x cannot by 2 nor -2

OpenStudy (anonymous):

right I had that already

OpenStudy (anonymous):

so it should just be what we found, but I checked and its not

OpenStudy (science0229):

So, the only possible x-values are sqrt(4/3) and -sqrt(4/3) Now, the slope HAS to change its sign from the inflection point's left to right.

OpenStudy (anonymous):

funny thing you say that. I should have known. the function is undefine at -2,2, right? So it does not change signs even thogh the concavity does.. so there are none...lol

OpenStudy (anonymous):

in the original function.ughhh, these inflection points always trip me up!

OpenStudy (anonymous):

Thanks for all of the help, people!

OpenStudy (anonymous):

how about help determining its domain and range?

OpenStudy (science0229):

Domain=All real numbers except 2 and -2